# LW-FedSSL: Resource-efficient Layer-wise Federated Self-supervised Learning

### v1

@@ 4.2 LW-FedSSL
where $z=(z^{+},\{z_{j}^{-}\}^{K}_{j=1})$. This loss can be integrated into local training as an auxiliary loss for regularization, with a weight term $\alpha$ to control its contribution. The representation alignment mechanism also helps local models stay close to the global model that has gone through training at the server.
Integrating the aforementioned mechanisms improves the effectiveness of layer-wise training, enabling LW-FedSSL to perform comparably with end-to-end training, all the while substantially reducing resource demands on FL clients. The details of LW-FedSSL are presented in Algorithms 1 and 2. Algorithm 1 outlines the server-side execution, which handles gradual increase in the model depth based on stage, server-side calibration mechanism, and aggregation. Each client executes Algorithm 2, performing local SSL training with representation alignment for regularization.
@@
Table 2: Linear evaluation and fine-tuning performance of different approaches. We conduct the experiments three times with different seeds. Three best values, except for the centralized MoCo v3, are marked in bold. We compare the memory, FLOPs, and communication costs (download + upload) for a local training process. For FedSimCLR, we use a batch size of 512.
@@
Figure 3: Computation and communication resources required for a local training process.
4.3 Prog-FedSSL
Inspired by Wang et al. Wang et al. (2022), we explore progressive training for FedSSL, denoted as Prog-FedSSL. Similar to LW-FedMoCoV3, progressive training also starts with an empty encoder $F$, appending a new layer $L_{s}$ to $F$ at each stage $s$. However, in contrast to updating only a single layer $L_{s}$, progressive training updates all existing layers (i.e., $L_{1}$ to $L_{s}$) during each stage $s$. This distinction results in higher computation and communication costs for Prog-FedSSL than layer-wise training approaches. Nonetheless, when compared with end-to-end training, which requires clients to train a full model with $S$ layers (i.e., $L_{1}$ to $L_{S}$) at each round, Prog-FedSSL enables clients to train only a sub-model with $s$ layers $(s\leq S)$. Moreover, clients only need to exchange the sub-model with the server. In essence, Prog-FedSSL saves both computation and communication costs for $(S-s)$ layers at each stage $s$ compared to end-to-end training. As we will demonstrate in Section 5.2, progressive training outperforms both layer-wise and end-to-end training approaches. Nevertheless, it is important to note that the memory requirement of progressive training gradually increases as the stages progress, becoming equivalent to that of end-to-end training at the final stage (i.e., when $s=S$). This may limit low-memory devices from participating in the FL process, whereas the memory requirement of layer-wise training is significantly lower due to training a single layer $L_{s}$ at a time.
5 Experiment
5.1 Experimental setup
@@
For performance comparison, we primarily use linear evaluation Kolesnikov et al. (2019); Bachman et al. (2019); Chen et al. (2020) in which a linear classifier is trained on top of the frozen encoder $F$. The classifier is trained for 20 epochs using a batch size of 256. We use the AdamW optimizer with a base learning rate of 3e-2 and a weight decay of 1e-5. The learning rate decays with a cosine schedule. We use the RandAugment function available in the PyTorch Paszke et al. (2019) framework for augmentation. For fine-tuning, we set the base learning rate to 1e-3 and the number of epochs to 40 with a warmup period of 10 epochs. Other settings are similar to those of linear evaluation.
@@
Figure 4: Linear evaluation performance of LW-FedSSL, the baseline LW-FedMoCoV3, and its two variants incorporating either server-side calibration or representation alignment. Weight transfer between blocks is not performed for these two variants, as they demonstrate better performance without it.
5.2 Performance and Resource Comparison
@@
Table 2 compares the performance of different approaches. In terms of linear evaluation, LW-FedSSL demonstrates comparable performance to the end-to-end training of FedMoCoV3 and significantly outperforms the purely layer-wise training of LW-FedMoCoV3. LW-FedSSL also outperforms or holds a competitive position against FedMoCoV3 in fine-tuning. The progressive training approach, Prog-FedSSL, outperforms all other approaches in both linear evaluation and fine-tuning performance. Interestingly, the fine-tuning performance of FedMoCoV3 is even lower than that of LW-FedMoCoV3 in some cases. Please note that we apply the same fine-tuning settings, which may not be optimal for all approaches.
@@
In Figure 3, we present resource requirements for a single client. Figure 3a measures the memory usage at each local epoch, showing that the memory requirement of LW-FedSSL is $\times3$ lower than that of FedMoCoV3 at peak usage. The memory usage of Prog-FedSSL gradually increases and eventually matches that of FedMoCoV3 in the final stage. Figure 3b shows that LW-FedSSL also maintains a lower FLOPs requirement than FedMoCoV3, with 2.1×2.1$\times$2.1 × lower total FLOPs consumption. Additional memory and FLOPs requirements of LW-FedSSL over LW-FedMoCoV3 arise from the representation alignment mechanism, which requires inference on the global model to obtain global representations.
@@
In Figures 3c and 3d, we investigate the communication cost required for model exchange (download and upload) between the server and a client. Note that we only consider the encoder part for comparison, excluding the projection and prediction heads, which are common to all approaches with constant communication costs. In FedMoCoV3, a client needs to exchange the full encoder with the server at each communication round. Meanwhile, in LW-FedMoCoV3, a client only needs to exchange the active layer (i.e., $L_{s}$). Thus, both FedMoCoV3 and LW-FedMoCoV3 result in constant download and upload costs across different communication rounds. For Prog-FedSSL, both download and upload costs gradually increase as the number of active layers (i.e., $L_{1}$ to $L_{s}$) grows with the stage $s$. As for LW-FedSSL, a client needs to download layers $L_{1}$ to $L_{s}$ from the server, as these are updated through the server-side calibration mechanism. However, the client uploads a single layer $L_{s}$, as only it is updated by the local training process. Therefore, the download cost of LW-FedSSL increases with each step $s$, while the upload cost remains constant. Upon comparing the total communication cost over 180 rounds, LW-FedSSL is 1.8×1.8$\times$1.8 × cheaper in terms of download cost and 12×12$\times$12 × cheaper in terms of upload cost compared to FedMoCoV3.
5.3 Ablation Study
Figure 5: Impact of batch size on different approaches. (a) Linear evaluation results on the STL-10 dataset with different batch sizes. (b) Peak memory requirements for different batch sizes.
@@ 5.3 Ablation Study
In Figure 4, we compare LW-FedSSL with the baseline LW-FedMoCoV3 and its two variants, which solely incorporate either server-side calibration or representation alignment. Incorporating server-side calibration improves performance on the CIFAR-10 dataset (from which the auxiliary data is sampled), but leads to a decline on other datasets. This may result from training on auxiliary data, whose update direction conflicts with that of local training. Compared to the baseline, incorporating representation alignment obtains better performance on the STL-10 dataset, with comparable performance on the others. Meanwhile, LW-FedSSL, which incorporates both mechanisms, consistently outperforms the baseline across all datasets. One plausible explanation for this is that representation alignment encourages local representations to stay close to global representations during local training. This in turn helps preserve the knowledge acquired from local data during the server-side calibration process. Figure 4 suggests that instead of acting independently, server-side calibration and representation alignment are complementary for improving the performance of layer-wise training.
5.4 Batch Size
@@
SSL approaches tend to perform better with larger batch sizes due to a greater number of negative samples Chen et al. (2020). To investigate this, we set the local training batch size to 64, 128, 256, 512, and 1024. Figure 5a indicates that layer-wise training approaches, LW-FedMoCoV3 and LW-FedSSL, also tend to improve with increased batch size, similar to end-to-end SSL training. Figure 5b reports peak memory requirements for each setting. Given that Prog-FedSSL does not freeze the layers, its peak memory requirement (at the last stage) is equivalent to that of FedMoCoV3. Memory requirements for FedMoCoV3 and Prog-FedSSL sharply rise as the batch size grows, while those of LW-FedMoCoV3 and LW-FedSSL remain relatively flat.
(a) Linear evaluation of LW-FedMoCoV3.
@@
Figure 7: Linear evaluation results on the STL-10 dataset. (a) Impact of different levels of data heterogeneity. (b) Impact of different numbers of clients.
Figure 8: Learning rate strategies.
5.5 Data Heterogeneity
In this experiment, we use the Tiny ImageNet dataset to construct the local data of clients. We control the data heterogeneity based on data labels and the Dirichlet distribution Ferguson (1973). The concentration parameter $\beta$ in the Dirichlet distribution determines the strength of data heterogeneity, with a lower $\beta$ value indicating a higher degree of heterogeneity. The labels are used here to set up the clients’ data partitions and are not used during training (except for the supervised baseline). Figure 7a presents the linear evaluation results for different $\beta$ values. For the supervised baseline, an end-to-end classification model is trained with vanilla FL using labeled data, and then only the encoder part is kept for linear evaluation. Notably, performance of the supervised approach diminishes as the degree of data heterogeneity increases. In contrast, other approaches employing SSL are relatively more robust to different $\beta$ values. This observation conforms with prior studies Wang et al. (2023) that highlight the robustness of SSL to data heterogeneity in FL environments. Figure 7a indicates that layer-wise training also exhibits robustness to data heterogeneity inherent in end-to-end SSL training.
5.6 Number of Clients
To investigate the impact of the number of clients $N$, we conduct experiments with $N$ set to 5, 10, 30, and 45. For each setting, we uniformly partition the unlabeled data portion of the STL-10 dataset into $N$ local datasets. Therefore, the total number of data samples remains constant across all settings. Linear evaluation results with varying $N$ values are shown in Figure 7b. We can observe that all approaches, except LW-FedSSL, experience a drop in performance as the number of clients increases. Integration of representation alignment mechanism enables LW-FedSSL to maintain good performance even in scenarios with a large number of clients Li et al. (2021).
5.7 Learning Rate Strategy
@@
By default, we use cosine decay scheduling for the learning rate. We explore different learning rate strategies for LW-FedMoCoV3, LW-FedSSL, and Prog-FedSSL, including a fixed learning rate and cosine decay scheduling for each training stage, which we refer to as “cyclic”. Figure 8 illustrates how the learning rate changes at each local epoch for different strategies, and the results are presented in Figure 6. For LW-FedMoCoV3, using a fixed learning rate achieves the best performance, and also for both LW-FedSSL and Prog-FedSSL, simply using a fixed learning rate can be a decent choice. A fixed learning rate may allow layers in later stages to receive effective updates, whereas in cosine decay scheduling, later layers undergo minimal updates due to a small learning rate. While the cyclic learning rate pattern proves to work decently for Prog-FedSSL, it may not be suitable for LW-FedMoCoV3 and LW-FedSSL. As seen in Figure 8, the learning rate for cyclic strategy drops quickly in each stage when only a limited number of rounds are available. This can impact layer-wise training, where a layer receives training only in its corresponding stage. (For LW-FedSSL with a fixed learning rate, we do not use weight transfer, as it demonstrates improved performance.)
6 Conclusion

### v2

@@ IV-B LW-FedSSL
where $z=(z^{+},\{z_{j}^{-}\}^{K}_{j=1})$. This loss can be integrated into local training as an auxiliary loss for regularization, with a weight term $\alpha$ to control its contribution.
The server-side calibration mechanism trains all existing layers in the global model to facilitate smooth collaboration between them. However, during local training, updates only to the active layer may disrupt this cohesion. To mitigate this, representation alignment encourages the updates of the local model to remain closely aligned with the parameters of the global model. This prevents complete overwriting and helps preserve the layer cohesion established by server-side calibration, while also allowing local models to incorporate local information from their respective local datasets. Additionally, representation alignment prevents extensive divergence between different local models, enabling a more effective FL aggregation process. In Section V-C, we empirically show that server-side calibration and representation alignment mechanisms are complementary for improving the effectiveness of layer-wise training. Integrating these mechanisms enables LW-FedSSL to compete with end-to-end training, all the while substantially reducing resource demands on FL clients. The details of LW-FedSSL are presented in Algorithms 1 and 2. Algorithm 1 outlines the server-side execution, which handles gradual increase in the model depth based on stage, server-side calibration mechanism, and aggregation. Each client executes Algorithm 2, performing local SSL training with representation alignment for regularization.
TABLE III: Linear evaluation and fine-tuning performance of different approaches. We conduct the experiments three times with different seeds. We compare the memory, FLOPs, and communication costs (download + upload) for a local training process. FedMoCo, FedBYOL, and FedSimCLR correspond to MoCo v3 [11], BYOL [29], and SimCLR [1] integrated with FL. For FedSimCLR, we use a batch size of 512. Three best values, except for the centralized MoCo v3, are marked in bold.
@@
Figure 5: Computation and communication resources required for a local training process. (a) Memory requirement. (b) FLOPSs consumption. (c) Download size. (d) Upload size.
IV-C Prog-FedSSL
Inspired by ProgFed in [18], we also explore progressive training for FedSSL, denoted as Prog-FedSSL. Similar to layer-wise training, progressive training also starts with an empty encoder $F$, appending a new layer $L_{s}$ to $F$ at each stage $s$. However, in contrast to updating only a single layer $L_{s}$, progressive training updates all existing layers (i.e., $L_{1}$ to $L_{s}$) during each stage $s$. This distinction results in higher computation and communication costs for Prog-FedSSL than layer-wise training approaches. Nonetheless, when compared with end-to-end training, which requires clients to train a full model with $S$ layers (i.e., $L_{1}$ to $L_{S}$) at each round, Prog-FedSSL enables clients to train only a sub-model with $s$ layers $(s\leq S)$. Moreover, clients only need to exchange the sub-model with the server. In essence, Prog-FedSSL saves both computation and communication costs for $(S-s)$ layers at each stage $s$ compared to end-to-end training. As we will demonstrate in Section V-B, progressive training outperforms both layer-wise and end-to-end training approaches. Nevertheless, it is important to note that the memory requirement of progressive training gradually increases as the stages progress, becoming equivalent to that of end-to-end training at the final stage (i.e., when $s=S$). This may limit low-memory devices from participating in the FL process, whereas the memory requirement of layer-wise training is significantly lower due to training a single layer $L_{s}$ at a time. We compare the different characteristics of FedMoCo (end-to-end), LW-FedMoCo, LW-FedSSL, and Prog-FedSSL in Table II. Fig. 15 in Appendix A illustrates FedMoCo, LW-FedMoCo, and Prog-FedSSL at stage $s$.
(a) Linear evaluation.
(b) Peak memory requirement.
Figure 6: Impact of batch size on different approaches. (a) Linear evaluation results on the STL-10 dataset and (b) peak memory requirements for different batch sizes.
V Experiment
V-A Experimental setup
@@
Following common practices, we primarily use linear evaluation for performance comparison [36, 37, 1]. The projection head $H$ and prediction head $P$ are discarded, and a linear classifier is trained on top of the frozen encoder $F$ for 20 epochs using a batch size of 256. We use the AdamW optimizer with a base learning rate of 3e-2 and a weight decay of 1e-5. The learning rate decays with a cosine schedule. We use the RandAugment function available in the PyTorch [38] framework for augmentation. For fine-tuning, we set the base learning rate to 1e-3 and the number of epochs to 40 with a warmup period of 10 epochs. Other settings are similar to those of linear evaluation.
@@
Figure 7: Linear evaluation performance of LW-FedSSL, the baseline LW-FedMoCo, and its two variants incorporating either server-side calibration or representation alignment. Weight transfer between blocks is not performed for these two variants, as they demonstrate better performance without it.
V-B Performance and Resource Comparison
@@
Table III compares the performance of different approaches. In terms of linear evaluation, LW-FedSSL demonstrates comparable performance to the end-to-end training of FedMoCo and significantly outperforms the purely layer-wise training of LW-FedMoCo. LW-FedSSL also outperforms or holds a competitive position against FedMoCo in fine-tuning. The progressive training approach, Prog-FedSSL, outperforms all other approaches in both linear evaluation and fine-tuning performance. Interestingly, the fine-tuning performance of FedMoCo is even lower than that of LW-FedMoCo in some cases. Please note that we apply the same fine-tuning settings, which may not be optimal for all approaches.
@@
In Fig. 5, we present resource requirements for a single client. Fig. 5a measures the memory usage at each local epoch, showing that the memory requirement of LW-FedSSL is $\times3$ lower than that of FedMoCo at peak usage. The memory usage of Prog-FedSSL gradually increases and eventually matches that of FedMoCo in the final stage. Fig. 5b shows that LW-FedSSL also maintains a lower FLOPs requirement than FedMoCo, with 2.1×2.1$\times$2.1 × lower total FLOPs consumption. Additional memory and FLOPs requirements of LW-FedSSL over LW-FedMoCo arise from the representation alignment mechanism, which requires inference on the global model to obtain global representations.
@@
In Figures 5c and 5d, we investigate the communication cost required for model exchange (download and upload) between the server and a client. Note that we only consider the encoder part for comparison, excluding the projection and prediction heads, which are common to all approaches with constant communication costs. In FedMoCo, a client needs to exchange the full encoder with the server at each communication round. Meanwhile, in LW-FedMoCo, a client only needs to exchange the active layer (i.e., $L_{s}$). Thus, both FedMoCo and LW-FedMoCo result in constant download and upload costs across different communication rounds. For Prog-FedSSL, both download and upload costs gradually increase as the number of active layers (i.e., $L_{1}$ to $L_{s}$) grows with the stage $s$. As for LW-FedSSL, a client needs to download layers $L_{1}$ to $L_{s}$ from the server, as these are updated through the server-side calibration mechanism. However, the client uploads a single layer $L_{s}$, as only it is updated by the local training process. Therefore, the download cost of LW-FedSSL increases with each step $s$, while the upload cost remains constant. Upon comparing the total communication cost over 180 rounds, LW-FedSSL is 1.8×1.8$\times$1.8 × cheaper in terms of download cost and 12×12$\times$12 × cheaper in terms of upload cost compared to FedMoCo.
Figure 8: Different levels of data heterogeneity. Linear evaluation results are shown for the STL-10 dataset.
TABLE IV: Linear evaluation results of LW-FedSSL with different amounts of auxiliary data.
Ratio CIFAR-10 CIFAR-100 STL-10
Tiny
ImageNet
10% ($D^{g}$ only) 43.75 20.03 40.97 9.77
1% 47.80 23.69 43.65 12.76
10% 61.83 32.25 55.15 17.26
100% 79.63 47.09 69.52 26.75
V-C Ablation Study
@@
In Fig. 7, we compare LW-FedSSL with the baseline LW-FedMoCo and its two variants, which solely incorporate either server-side calibration or representation alignment. Incorporating server-side calibration improves performance on the CIFAR-10 dataset (from which the auxiliary data is sampled), but leads to a decline on other datasets. This decline may result from conflicting update directions between training on auxiliary data and the local training. Compared to the baseline, incorporating representation alignment obtains better performance on the STL-10 dataset, with comparable performance on the others. Meanwhile, LW-FedSSL, which incorporates both mechanisms, consistently outperforms the baseline across all datasets. The server-side calibration mechanism facilitates smooth collaboration between different layers of the model. Meanwhile, representation alignment encourages the weights of the local model to remain closely aligned with those of the global model. This helps maintain layer cohesion established by the server-side calibration mechanism during the local training process. Fig. 7 suggests that instead of acting independently, server-side calibration and representation alignment are complementary for improving the performance of layer-wise training.
V-D Amount of Auxiliary Data
The server-side calibration mechanism of LW-FedSSL relies on the auxiliary dataset $D^{g}$. We vary the amount of auxiliary data at the server by sampling 1%, 10%, and 100% of the CIFAR-10 training data. We also conduct a centralized end-to-end SSL training using only the auxiliary data at the sever. The results in Table IV indicate that the performance of LW-FedSSL improves with an increase in the amount of data.
V-E Batch Size
@@
SSL approaches tend to perform better with larger batch sizes due to a greater number of negative samples [1]. To investigate this, we set the local training batch size to 64, 128, 256, 512, and 1024. Fig. 6a indicates that layer-wise training approaches, LW-FedMoCo and LW-FedSSL, also tend to improve with increased batch size, similar to end-to-end SSL training. Fig. 6b reports peak memory requirements for each setting. Given that Prog-FedSSL does not freeze the layers, its peak memory requirement (at the last stage) is equivalent to that of FedMoCo. Memory requirements for FedMoCo and Prog-FedSSL sharply rise as the batch size grows, while those of LW-FedMoCo and LW-FedSSL remain relatively flat.
Figure 9: Total number of clients. Linear evaluation results are shown for the STL-10 dataset.
Figure 10: Number of participating clients in each FL round. The total number of clients is set to 45. Linear evaluation results are shown for the STL-10 dataset.
V-F Data Heterogeneity
Data heterogeneity is one of the main challenges in an FL environment. In this experiment, we use the Tiny ImageNet dataset to construct the local data of clients. We control the data heterogeneity based on data labels and the Dirichlet distribution [39]. The concentration parameter $\beta$ in the Dirichlet distribution determines the strength of data heterogeneity, with a lower $\beta$ value indicating a higher degree of heterogeneity. The labels are used here to set up the clients’ data partitions and are not used during training (except for the supervised baseline). Fig. 8 presents the linear evaluation results for different $\beta$ values. For the supervised baseline, an end-to-end classification model is trained with vanilla FL using labeled data, and then only the encoder part is kept for linear evaluation. Notably, performance of the supervised approach diminishes as the degree of data heterogeneity increases. In contrast, other approaches employing SSL are relatively more robust to different $\beta$ values. This observation conforms with prior studies [8] that highlight the robustness of SSL to data heterogeneity in FL environments. Fig. 8 indicates that layer-wise training also exhibits robustness to data heterogeneity inherent in end-to-end SSL training.
(a) Uniform allocation.
(b) Right-skewed allocation. (high to low)
(c) Left-skewed allocation. (low to high)
Figure 11: Number of rounds per stage.
(a) Linear evaluation of LW-FedMoCo.
(b) Linear evaluation of LW-FedSSL.
(c) Linear evaluation of Prog-FedSSL.
Figure 12: Impact of number of rounds per stage.
Figure 13: Learning rate strategies.
V-G Total Number of Clients
To investigate the impact of the number of clients $N$, we conduct experiments with $N$ set to 5, 10, 30, and 45. For each setting, we uniformly partition the unlabeled data portion of the STL-10 dataset into $N$ local datasets. Therefore, the total number of data samples remains constant across all settings. Linear evaluation results with varying $N$ values are shown in Fig. 9. We can observe that all approaches, except LW-FedSSL, experience a drop in performance as the number of clients increases. Integration of representation alignment mechanism enables LW-FedSSL to maintain good performance even in scenarios with a large number of clients [27].
V-H Number of Participating Clients
In FL environments, it is common for only a subset of clients to participate in training at any given time. Therefore, we experiment with the number of clients active for training at each FL round. We set the total number of clients to 45 and randomly sample 5, 10, 20, and 45 clients at each FL round. The linear evaluation results in Fig. 10 demonstrate that LW-FedSSL can maintain its performance even with a small proportion of participating clients in each FL round.
V-I Distribution of Number of Rounds Per Stage
In our experiments, the number of rounds for each layer-wise training stage is uniformly distributed. As shown in Fig. 11, we explore two alternative settings: one where a larger number of rounds is allocated to earlier stages, and the other where a larger number of rounds is allocated to later stages. We refer to these settings as “right-skewed allocation” and “left-skewed allocation”, respectively. The total number of rounds across all settings remains constant at 180. The results are shown in Fig. 12.
For LW-FedMoCo, the left-skewed setting results in the worst performance. In this setting, earlier layers receive updates for a small number of rounds, potentially lowering the quality of their output features. Even though later layers are trained for a large number of rounds, poor output features from the early layers may degrade the output of subsequent layers. The other two settings achieve comparable performance. As for LW-FedSSL, it allows early layers to continuously receive updates through the server-side calibration mechanism, which improves the performance of the left-skewed setting in most cases.
In the case of Prog-FedSSL, none of the layers from prior stages are frozen, allowing a layer $L_{s}$ to receive training for all the rounds allocated through stages $s$ to $S$. This means that in a left-skewed allocation scenario, if we assign 10 rounds to stage $S-1$ and 50 rounds to stage $S$, then the layer $L_{S-1}$ undergoes training for a total of 60 rounds (since it is not frozen), while $L_{S}$ is trained for 50 rounds. However, if we reverse the scenario in the right-skewed allocation, i.e., assign 50 rounds to stage $S-1$ and 10 rounds to stage $S$, the layer $L_{S-1}$ still undergoes training for 60 rounds, while $L_{S}$ is trained only for 10 rounds. As a result, the left-skewed setting achieves better performance than the right-skewed setting; however, it would also consume more computation resources. The same case also applies to the server-side calibration mechanism of LW-FedSSL.
(a) Linear evaluation of LW-FedMoCo.
@@ V-I Distribution of Number of Rounds Per Stage
V-J Learning Rate Strategy
@@
By default, we use cosine decay scheduling for the learning rate. We explore different learning rate strategies for LW-FedMoCo, LW-FedSSL, and Prog-FedSSL, including a fixed learning rate and cosine decay scheduling for each training stage, which we refer to as “cyclic”. Fig. 13 illustrates how the learning rate changes at each local epoch for different strategies, and the results are presented in Fig. 14. For LW-FedMoCo, using a fixed learning rate achieves the best performance, and also for both LW-FedSSL and Prog-FedSSL, simply using a fixed learning rate can be a decent choice. A fixed learning rate may allow layers in later stages to receive effective updates, whereas in cosine decay scheduling, later layers undergo minimal updates due to a small learning rate. While the cyclic learning rate pattern proves to work decently for Prog-FedSSL, it may not be suitable for LW-FedMoCo and LW-FedSSL. As seen in Fig. 13, the learning rate for cyclic strategy drops quickly in each stage when only a limited number of rounds are available. This can impact layer-wise training, where a layer receives training only in its corresponding stage. (For LW-FedSSL with a fixed learning rate, we do not use weight transfer, as it demonstrates improved performance.)
VI Conclusion